Moisa Evans is a maitre d' at Red Rock Club. On September 6, his gross pay was $900 (three days working, one paid vacation day, and one paid sick day). He also reported to his employer tips of $860 for the previous month (applicable taxes to be deducted out of this pay). Evans belongs to the company's 401(k) plan and has 5% of his gross pay ($900) deducted each week (salary reduction). Red Rock Club also provides a matching contribution ($45) into the plan for Evans. Compute the following amounts:

a. Deduction for OASDI tax $
b. Deduction for HI tax $
Answer #1

a.

deduction for OASDI tax

$109.12

b.

deduction for HI tax

$25.52

Moisa Evans's gross income for OASDI tax and HI tax = Gross pay + Employer tips = 900+860 = $1760

(contribution to 401(k) is not considered for OASDI and HI taxes)

Therefore,

Deduction for OASDI tax = gross income *6.2% =1760*6.2% = $109.12

Deduction for HI tax = gross income *1.45% = 1760*1.45% = $25.52